OpenStudy (anonymous):

Tell whether the ordered pair (3, -1) is a solution of the given system. {x-2y=5 2x-y=7 Answers: Yes, No

OpenStudy (amistre64):

when (x=3 and y=-1) do we get true statements or false statements?

OpenStudy (amistre64):

x-2y=5 3 -1 -------- 3+2 =5 , yes or no? 2x-y=7 3 -1 ------- 6+1 = 7 ; yes or no? if both are yes, then its a solution to both; if one or none or no; then its a bad solution